OpenStudy (anonymous):

Solve: Limx-->0+ sinx/(x^2) Limx-->0- sinx/(x^2) To prove Lim x-->0 sinx/(x^2) DNE without using L'Hopital's Rule

OpenStudy (anonymous):

How would I do this: 1. Without using squeeze theorem 2. With squeeze theorem?

OpenStudy (anonymous):

without: plug 0+ (0.0000001) and 0- (0.0000001) in your calculator to prove that lim 0+ DNE 0- therefore limit DNE. with: \[-1 \le sinx \le 1\]\[-\frac{ 1 }{ x^2 }\le\frac{ sinx }{ x^2 }\le \frac{ 1 }{ x^2 }\]\[lim_{x->0}(-\frac{ 1 }{ x^2 })\le\frac{ sinx }{ x^2 }\le lim_{x->0}(\frac{ 1 }{ x^2 })\] \[- \infty \le \lim_{x->0} \frac{sinx}{x^2} \le \infty\] since left and right aren't the same. we can not deduce the limit of sinx/x^2 by squeeze theorem since it DNE

OpenStudy (anonymous):

0- = (-0.0000001)

OpenStudy (anonymous):

the 3rd line of squeeze theorem should be lim of sinx/x^2 as well. we took the limit of all 3 terms at the same time for this to be true. feel free to ask questions ^_^

OpenStudy (anonymous):

I get it! Thank you! How would you solve this without squeeze theorem? I'm actually trying to find the vertical asymptote using limits!

OpenStudy (anonymous):

by graphing. should look like this at zero:  or plug in 0- and 0+ on your calculator. 0- means 0 coming form the left. the closest number to zero from the left is -0.0000000000000.........1 and 0+ is coming from the right of zero. so you can plug in +0.0000000....1 with you calculator, not that many 0s are necessary to estimate if you get the same result or not